A galvanometer of resistor 10 ohms gives a full scale deflection when a current of 50amphere flow through it.

How can it be converted to an ammeter to measure a current of 5Amphere?

How can it be converted to a voltmeter to measure a voltage of 3volts?


1
Expert's answer
2021-11-18T12:21:59-0500

(a) Ammeter


A galvanometer can function as an ammeter when it is placed in parallel with a small resistance "R" .


"IR=I_Gr\to R=\frac{I_G}{I}r=\frac{0.000005}{5-0.000005}\cdot 10=1\cdot10^{-5} \ (\Omega)"


(b) Voltmeter


A galvanometer can function as a voltmeter when it is connected in series with a large resistance "R" .

"R_{total}=R+r=V/I=3/(50\cdot10^{-6})=60000\ (\Omega)"


"R=R_{total}-r=60000-10=59990\ (\Omega)"